To stitch a shirt, 2 m 15 cm cloth is needed. Out of 40 m cloth, how many shirts can be stitched and how much cloth will remain?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
Total length of the cloth = 40 m = 40×100 cm = 4000 cm
Cloth required to stitch one shirt = 2 m 15 cm = 2×100 + 15 cm = 215 cm
Number of shirts that can be stitched out of 4000 cm = 4000 / 215 = 18 shirts
Hence 18 shirts can be stitched out of 40 m and 1m 30 cm of cloth is left out

Thanks